OpenStudy (anonymous):

@Studygirl14 Please verify my answer for me .

OpenStudy (anonymous):

That is the cylinder..

OpenStudy (studygurl14):

Everything looks great. Awesome job. :)

OpenStudy (anonymous):

Really? Awesome thank you. :)

OpenStudy (studygurl14):

Yep, really. See, it isn't that hard. :)

OpenStudy (anonymous):

No just memorizing formulas

OpenStudy (studygurl14):

Wait, hold up

OpenStudy (studygurl14):

Is this surface area of lateral area?

OpenStudy (studygurl14):

or*

OpenStudy (anonymous):

Wait I got it wrong I think...

OpenStudy (anonymous):

No I don't think so. Just surface area no lateral.

OpenStudy (studygurl14):

You found the lateral area, not the surface area...i think that problem called for surface area

OpenStudy (studygurl14):

okay, so then sorry. incorrect

OpenStudy (anonymous):

yeah damn.

OpenStudy (studygurl14):

Find the area of one of the circles, multiply it by 2, then add it to the answer you got for lateral area. make sure to leave pi

OpenStudy (studygurl14):

Do you know how to find the area of a circle?

OpenStudy (anonymous):

Um yes pi * r^2 ?

OpenStudy (anonymous):

so 3.14 * 16?

OpenStudy (studygurl14):

don't use the 3.14. leave it as pi. And yes 16

OpenStudy (studygurl14):

because there are two circles that are identical you multiply \(\large 16\pi\) by 2. What do you get?

OpenStudy (anonymous):

wait so my answer for the area of circle is 50.265?

OpenStudy (studygurl14):

no....

OpenStudy (studygurl14):

what is 16 x 2?

OpenStudy (anonymous):

okay so 32.

OpenStudy (studygurl14):

right. So you get \(\large 32\pi\) And remember you have \(\large 48\pi\) from before. Add them together, and leave pi as pi. Treat pi as a varable

OpenStudy (studygurl14):

So... \(\large 32\pi + 48\pi=?\)

OpenStudy (anonymous):

80 pi.

OpenStudy (studygurl14):

yep

OpenStudy (anonymous):

alright thank you

OpenStudy (studygurl14):

you're welcome